# Enhancing RAG Efficiency with Adaptive Context Compression

### v1

@@ Abstract
Enhancing RAG Efficiency with Adaptive Context Compression
@@ 1 Introduction
We propose an adaptive context compression framework for RAG that dynamically adjusts the number of compressed embeddings during inference. The framework decouples offline compression (fixed-rate hierarchical embeddings) from online selection (dynamic embedding feeding, halted once sufficient context is accumulated), mimicking human selective reading. We establish a unified benchmark (Wikipedia corpus + five QA datasets) to ensure fair evaluation. Our method outperforms other compression techniques in effectiveness and efficiency, matches or exceeds standard RAG accuracy on four datasets with 4×4$\times$4 × faster inference, and demonstrates significant potential for adaptive compression via compressor analysis.
@@ 3 Adaptive Context Compression for RAG
Figure 2: Overview of Adaptive Context Compression for RAG pipeline. The workflow is as follow: Context Retrieval $\xrightarrow{}$ Hierarchical Compression $\xrightarrow{}$ Adaptive Context Selection $\xrightarrow{}$ Response Generation.
@@ 3.1.3 Multi-granularity Compression
In addition to encoding useful information, the compressor must also learn hierarchical information organization. We define granularity $b$ as the starting position of an embedding sequence. The compressed sequence should exhibit varying information densities across subsequences truncated at different granularity levels. To achieve multi-granular compression, we optimize across multiple granularities during both pretraining and fine-tuning. Formally, given a training granularity sequence =$\mathcal{B}=\{b_{1},\ldots,b_{k}\}$, the training objective combines reconstruction errors at all granularities:
@@ 3.2 Adaptive Selector
The adaptive selector $\mathcal{S}$ dynamically controls decoder input by progressively accumulating and verifying context embeddings during inference. The query is encoded into a query embedding sequence $E_{q}=\mathcal{D}_{\text{embed}}(q)$, while the context embedding sequence is obtained by concatenating all compressed document sequences in retrieval order: $E_{c}=\mathrm{Concat}(E_{1},\ldots,E_{k})$. The selector iterates through granularity levels from the inference granularity sequence =$\mathcal{B}=\{b_{1},\ldots,b_{k}\}$. At each step $t$:
@@ 3.2.2 Reinforcement Learning
Action: $\quad a_{t}=\begin{cases}1&\text{sufficient}\\ 0&\text{continue}\end{cases}$
@@ 4.3 Evaluation Metrics
AutoCompressor ×$\times 4$ 24.13 1007 14.81 1242 28.15 1015 23.49 1085 15.79 1057 21.27 1081
COCOM ×$\times 4$ 26.73 1007 18.83 1242 23.92 1015 31.56 1085 25.60 1057 25.33 1081
@@
xRAG ×$\times 128$ 5.93 269 8.61 485 12.60 265 17.00 324 7.77 301 10.38 328
ICAE ×$\times 128$ 23.30 269 18.11 485 23.57 265 30.55 324 18.47 301 22.8 328
×$\times 16$ 27.04 440 19.23 667 23.97 436 31.41 502 23.95 476 25.12 504
×$\times 4$ 27.53 1007 19.63 1242 24.51 1015 31.41 1085 27.58 1057 26.13 1081
@@ 5.1 Main Results
The main results for ACC-RAG are presented in Table 1. Our ACC-RAG method, trained and inferenced with a granularity sequence of $[1,32]$, balances effectiveness and efficiency. Detailed results for other ACC-RAG configurations are in Section 5.4. The experimental results demonstrate that ACC-RAG significantly (Paired t-test, $p<0.05$) outperforms all compression baselines in M score while maintaining superior efficiency. Notably, under our unified benchmark, most methods (e.g., xRAG, COCOM) exhibit lower effectiveness compared to their original reports, primarily due to their reliance on larger-scale training data and task-specific tuning. In contrast, ACC-RAG achieves a balanced trade-off between effectiveness and efficiency: compared to direct generation, it significantly improves M scores on NQ, TriviaQA, and SQuAD (p<0.05) with minimal drop on WQ and TREC (p>0.05). Meanwhile, it matches or exceeds vanilla RAG’s accuracy on four datasets while reducing FTIT by over 4×4$\times$4 ×.

### v2

# Dynamic Context Compression for Efficient RAG

@@ Abstract
Dynamic Context Compression for Efficient RAG
@@ 1 Introduction
We propose an adaptive context compression framework for RAG that dynamically adjusts the number of compressed embeddings during inference. The framework decouples offline compression (fixed-rate hierarchical embeddings) from online selection (dynamic embedding feeding, halted once sufficient context is accumulated), mimicking human selective reading. We establish a unified benchmark (Wikipedia corpus + five QA datasets) to ensure fair evaluation. Our method outperforms other compression techniques in effectiveness and efficiency, matches or exceeds standard RAG accuracy on four datasets with $4\times$ faster inference, and demonstrates significant potential for adaptive compression via compressor analysis.
@@ 3 Adaptive Context Compression for RAG
[…]
@@ 3.1.3 Multi-granularity Compression
In addition to encoding useful information, the compressor must also learn hierarchical information organization. We define granularity $b$ as the starting position of an embedding sequence. The compressed sequence should exhibit varying information densities across subsequences truncated at different granularity levels. To achieve multi-granular compression, we optimize across multiple granularities during both pretraining and fine-tuning. Formally, given a training granularity sequence $\mathcal{B}=\{b_{1},\ldots,b_{k}\}$, the training objective combines reconstruction errors at all granularities:
@@ 3.2 Adaptive Selector
The adaptive selector $\mathcal{S}$ dynamically controls decoder input by progressively accumulating and verifying context embeddings during inference. The query is encoded into a query embedding sequence $E_{q}=\mathcal{D}_{\text{embed}}(q)$, while the context embedding sequence is obtained by concatenating all compressed document sequences in retrieval order: $E_{c}=\mathrm{Concat}(E_{1},\ldots,E_{k})$. The selector iterates through granularity levels from the inference granularity sequence $\mathcal{B}=\{b_{1},\ldots,b_{k}\}$. At each step $t$:
@@ 3.2.2 Reinforcement Learning
[…]
@@ 4.3 Evaluation Metrics
AutoCompressor $\times 4$ 24.13 1007 14.81 1242 28.15 1015 23.49 1085 15.79 1057 21.27 1081
COCOM $\times 4$ 26.73 1007 18.83 1242 23.92 1015 31.56 1085 25.60 1057 25.33 1081
@@
xRAG $\times 128$ 5.93 269 8.61 485 12.60 265 17.00 324 7.77 301 10.38 328
ICAE $\times 128$ 23.30 269 18.11 485 23.57 265 30.55 324 18.47 301 22.8 328
$\times 16$ 27.04 440 19.23 667 23.97 436 31.41 502 23.95 476 25.12 504
$\times 4$ 27.53 1007 19.63 1242 24.51 1015 31.41 1085 27.58 1057 26.13 1081
@@ 5.1 Main Results
The main results for ACC-RAG are presented in Table 1. Our ACC-RAG method, trained and inferenced with a granularity sequence of $[1,32]$, balances effectiveness and efficiency. Detailed results for other ACC-RAG configurations are in Section 5.4. The experimental results demonstrate that ACC-RAG significantly (Paired t-test, $p<0.05$) outperforms all compression baselines in M score while maintaining superior efficiency. Notably, under our unified benchmark, most methods (e.g., xRAG, COCOM) exhibit lower effectiveness compared to their original reports, primarily due to their reliance on larger-scale training data and task-specific tuning. In contrast, ACC-RAG achieves a balanced trade-off between effectiveness and efficiency: compared to direct generation, it significantly improves M scores on NQ, TriviaQA, and SQuAD (p<0.05) with minimal drop on WQ and TREC (p>0.05). Meanwhile, it matches or exceeds vanilla RAG’s accuracy on four datasets while reducing FTIT by over $4\times$.